A rectangular parcel of land is 70ft longer than it is wide. Each diagonal between opposite corbers is 130ft. What are the demensions of the parcel?
Answer by checkley71(8403)

GIVEN THAT THE DIAGONALS ARE EACH(=) THEN THESE DIAGONALS FOR A RIGHT TRIANGLE.
THUS:
X^2+(X+70)^2=130^2
X^2+X^2+140X+4900=16900
2X^2+140X+4900-16900=0
2X^2+140X-12000=0
2(X^2+70-6000)=0
2(X-50)(X+120)=0
X-50=0
X=50 FEET IS THE ANSWER FOR THE WIDTH.
50+70=120 FEET FOR THE LENGTH.
PROOF:
50^2+120^2=130^2
2500+14400=16900
16900=16900